# Advancing Strategic Planning and Dynamic Control of Complex Projects

### v1

@@ Abstract
Strategic project planning and dynamic control is essential for ensuring complex projects to be executed best-fit for common purpose. When best-for-project strategies are no longer conceivable for humans, there is a need for effective and efficient computer-aided decision support. To this end, standard simulation-driven evaluation and a-posteriori decision-making is often used instead of a combined simulation and optimisation approach that a-priori integrates the stakeholders goal orientation and risk management into a best-fit design solution. However, recently developed state-of-the-art planning and control methodologies that do already incorporate this integration still lack a stochastic representation and/or an associative multi-objective approach. This paper presents a new project management methodology, called Odycon (Open design and dynamic control), which dissolves the aforementioned shortcomings. To enable this, a generic mathematical statement is first formulated for project planning and dynamic control uniting technical (logistical) capabilities, human goal-oriented behaviour, and stakeholders interests. Then, both Monte-Carlo simulation (MCS) and the Integrative Maximisation of Aggregated Preferences (IMAP) optimisation method are combined, resulting in a best-fit strategic planning and dynamic control methodology. Odycon’s use and added value are demonstrated for two applications: (1) a pure strategic planning of an offshore wind installation project, and (2) a pure dynamic control of a highway infrastructure construction project. Both applications demonstrate Odycon’s advances in concurrent and associative design and decision-making, offering best-fit for common purpose synthesis for different complex project phases.
@@
Project management , Project scheduling , Project control , Risk management , Multi-objective optimisation , Monte Carlo simulation , Preference function modelling
journal: arXiv
@@ 1 Introduction
Successful project delivery is traditionally attempted through project planning and control, answering what is going to be done when and how much does it cost. Planning is referred to the process of deciding how to execute a project, while project control reflects the process of ensuring execution as planned. Establishing a robust and reliable project schedule is fundamental for preliminary cost and time estimates, as well as for effective and efficient project management (Smith et al., 2014). However, with increasing size and degree of uncertainty, initial estimates of costs and timelines tend to be fraught with inherent risks and uncertainty that cannot be completely eliminated. These uncertainties often results in discrepancies between early predictions and stakeholder objectives, complicating the decision-making process and may result in unmet expectations. Initial project planning and execution control is often approached by individual optimisation or simulation methods, resulting in a one-sided point of view and neglecting the multifaceted nature of projects where the technical capabilities, the human goal-oriented behaviour, and the stakeholders desirability are interconnected. It requires a systems-oriented approach to acknowledge these complexities and recognise the need for stakeholder-oriented decision-making to answer what should be accomplished why and how (Reschke and Schelle, 2013).
The role of well-informed human decision-making is crucial in project management, where decision-makers must understand the rationale behind their chosen project planning or control strategies towards risk management, maximum effectiveness, and ultimately stakeholder satisfaction. When evaluating the overall project’s success, it becomes clear that factors beyond just cost and timely completion play a significant role (Van Gunsteren, 2011). Trade-offs between quality, costs, and schedule are a constant reality in construction projects and are typically determined by the involved stakeholders. These trade-offs are reflected by the socio-technical interplay between the stakeholders preference (human domain) and the capability of the project (technical domain), see Van Heukelum et al. (2024). However, the traditional best practice recommends solely adhering to the contract and making a plan without considering the actual project reality, often following the rule to make a plan and stick to it. This practice can be conceived as ’pseudo-reality’, as project success is created by the dynamic adaptation to issues that arise (most often not defined in a contract) and dynamic interaction between stakeholders towards associative solutions. Despite this reality, the traditional managerial focus in construction projects often leans heavily towards costs and schedule and their individual interest, thus neglecting the interplay with stakeholder-oriented concurrent objectives in the process. When stakeholders are willing to go beyond their pure self-interest, joint project success can be enhanced. Stakeholders do have to relinquish their identified individual goals, ultimately enabling a synthesis solution that best fits the common purpose and outperforms single-sided project or even compromise solutions (Glasl, 1998; Scharmer, 2016; Van den Doel and Van Velthoven, 1993; Wolfert, 2023).
Traditional planning & control methods lack to explicitly express the relationship and influence of all participating stakeholder interests towards the decision-making process in project or operations management (Blanchard and Fabrycky, 2011; Del Pico, 2023; Slack et al., 2010). Usually (both in science and in industry), these methods are simulation and analysis oriented (a-posteriori) and do not take an (a-priori) design approach to effectively and efficiently generate a best-fit management solution. However, state-of-the-art methods that are optimisation-based often do not reflect a best-fit for common purpose consideration. These are most often being limited to a single objective optimisation (SOO) approach, relying on monetisation to reflect multiple objectives and neglecting the complexity of stakeholder satisfactory project management. While the benefit of multi-objective optimisation (MOO) in project management has been acknowledged (Guo and Zhang, 2022), there has been no methods developed towards adapting MOO in strategic project planning and dynamic project control integrating probabilistic risk management (Kammouh et al., 2022). Well-informed decision-making is facilitated by removing overconfidence bias through the use of structured and data-driven decision-making processes (Kahneman, 2011). Considering the inclusion of the goal-orientation of all involved stakeholders and their desired expectations, the increasing complexity of construction projects necessitates effective and efficient computer-aided decision support for stakeholder-oriented project management.
To address the lack of consideration for human control and adaptive behaviour in standard probabilistic scheduling methods using Monte-Carlo simulation (MCS) (Del Pico, 2023), research by Kammouh et al. (2022, 2021) has proposed the mitigation controller (MitC), a current state-of-the art project control software. It automates the project manager’s selective nature to only apply the most effective and efficient control measures on-the-run, based on the unique circumstances of each possible project realisation. The MitC, is designed as a SOO control tool, optimising for cost-effectiveness given a project delivery date (target constraint). As a result, there is no understanding of how the most optimal set of control measures impacts other project objectives from relevant decision-makers. While other objectives can be assessed independently, there is no guarantee that an optimal project planning and control strategy is derived that represents a best-fit for common purpose synthesis. Moreover, with increasing project complexity, the impact of individual planning & control variables related to all stakeholder objectives becomes cumbersome to be evaluated in a manual trial-and-error approach.
Associative project planning and control enabling a stakeholder-oriented behaviour can be introduced by utilising the advances developed in the Open Design Systems (Odesys) methodology (Wolfert, 2023), that enables the association of the individual stakeholder interests (desirability) into the technical design optimisation (capability) using integrative maximisation of aggregated preferences (IMAP). This Odesys/IMAP approach proves to be an effective methodology to arrive at best-fit for common purpose design solutions where all stakeholder interest are accounted for and translated to a common preference domain to find an overall group optimum. The introduced IMAP method enables iterative group design and decision-making in a deterministic manner. However, to effectively utilise these advancements for project planning and control, it is crucial to incorporate traditional probabilistic project planning and control methods that can precisely capture and represent the project’s schedule capabilities, such as the Program Evaluation and Review Technique (PERT). Additionally, these tools should employ stochastic simulation methods to reflect the inherent uncertainties and unpredictability encountered both in the planning and execution phases.
Utilising and combining these advances in simulation and optimisation can notably reduce inefficient and misinformed decision-making in project planning and control, particularly with increasing size of projects, where it becomes virtually impossible to account for all possible scenarios. In this context, the focus of this paper is to introduce a methodology that integrates (1) the physical and logistical capability of the project (’network performance domain’), with (2) the human goal-oriented behaviour (’decision-maker objective domain’) and extends it towards (3) the desirability of the project stakeholders (’associative preference domain’) enabling best-fit for common purpose project management. This is required as the current state-of-the-art methodologies for project planning or control exhibit the above-mentioned shortcomings that make them unsuitable for delivering the necessary integrative & associative behaviour to successfully execute the project so that it best meets all stakeholder interests. To achieve such a best-for-project fit, the project management practice PII introduced by Van Gunsteren (2011) and based on Schön and Argyris (1996) can be further improved to a PIII level. This advancement goes beyond the integrative double-loop learning approach, which involves making decisions for the stakeholder or project client. Instead, PIII emphasizes open-loop learning by working closely in association with all project stakeholders to achieve the best possible overall project result. A step from stakeholder oriented, but static ’self’-planning and control, to pure dynamic ’group’-planning and organisation where an integrative maximization of associated preferences is achieved Wolfert (2023).
Based upon the identified development gap and utilisation needs, this paper presents a new integrative stochastic simulation and multi-objective planning & control optimisation methodology, named Odycon (an acronym of open design and dynamic control). Odycon enables best-fit for common purpose generation of both strategic planning and/or dynamic control applications. Odycon is a decision-support methodology that automates the selection of planning and control variables, taking into consideration multiple stakeholder objectives and constraints. Building upon the MitC introduced by Kammouh et al. (2021) and the IMAP optimisation method introduced by Wolfert (2023) and Van Heukelum et al. (2024), this paper aims to advance open and dynamic project planning and control of complex construction projects, fostering multi-objective stakeholder-oriented decision-making. The following advances are presented:
(1) A general framework for an integrated strategic planning and dynamic control problem, which combines probabilistic project schedule simulations with a multi-objective design optimisation. This allows for dynamic adaptation to project-specific needs and stakeholder objectives, enhancing associative decision-making in complex construction projects.
(2) A direct representation between the project feasibility, project management objectives, and stakeholder decisions, achieving a unified integration of network performance, stakeholder objectives, and associative decision-making.
(3) A pure preference-based aggregation for multi-objective decision-making for project management, which translates various objectives and resources into a common preference domain to achieve an optimised group synthesis. This resolves the need for monetisation of non-cost-related objectives and does not go for the lowest cost but for the highest value.
@@
This paper first introduces the general mathematical statement of the multi-objective project planning & control optimisation. Following this, the underlying framework and concept of the stochastic simulation & optimisation methodology are described, with a detailed elaboration on the design of the method, in which the IMAP optimisation is implemented. The methodology is then demonstrated using a pure strategic planning example (no control variables) for an offshore wind installation project, where a service provider and a marine contractor are the key decision-makers. In contrast to the example focusing on pure strategic planning, the paper also demonstrates a pure dynamic control example (no planning variables) for a highway infrastructure project. This example includes associative control by both a contractor and a highway agency, showcasing the application of the methodology for two distinct project management phases. Finally, recommendations for further development are provided, and the added value of the newly developed stochastic simulation and optimisation methodology is concluded.
@@ 2 Mathematical formulation of the multi-objective optimisation
    $P_{k,i}(O_{i}(\mathcal{N}(\mathbf{x},\mathbf{y}),F_{1,2,...,J}(\mathbf{x},\mathbf{y}))$: Preference functions that describe the preference stakeholder $k$ has towards objective function $i$, which are functions of different performance functions and dependent on planning and/or control variables and physical variables ( $i\leq k$ and $K$ is the maximum number of stakeholders).
@@
    $\mathcal{N}(\mathbf{x},\mathbf{y})$: Performance function that describes the project network, depending on planning and/or control variables $\mathbf{x}$ and physical variables $\mathbf{y}$. Here $\mathcal{N}$ is a logical representation of a physical project planning network, composed of nodes (activities) and edges (logical links).
@@
The preference aggregation is performed using the principles of the PFM theory (Barzilai, 2022). It is an integral part of the integrative preference maximisation method (IMAP), see Wolfert (2023). To arrive at an aggregated preference for individual objectives, the IMAP method employs the mathematical operator, A, which determines the aggregated preference score using an algorithm called the A-fine Aggregator (see A).
@@
The preference functions describe the relation of a stakeholders satisfaction towards a certain objective on a defined scale. The elicitation of the preference functions and associated individual weights is essential to reflect the relationship of the stakeholders interest towards a certain objective. As an initial estimation for both the preference functions input and local and global the weights the Conjoint Analysis (CA) method can be used. For more information regarding the process of preference elicitation, the reader is referred to Arkesteijn et al. (2017). As mentioned in Van Heukelum et al. (2024) objective $O_{i}$ can be associated with multiple stakeholder preference functions $P_{k,i}$ (as $k\geq i$). Therefore, it is not required that a stakeholder expresses a preference for all objectives.
@@
In the context of a pure project planning optimisation (no control measures), the variables $x_{n}$ are defined as stated above, allowing to take any form (e.g. integer, float, binary, etc.). However, for a pure project control optimisation (no planning variables), the variables $\mathbf{x}$ are redefined as $\mathbf{x}=\mathbf{a}\mathbf{c_{i}}$, that is the product of the allocation of a control variables vector $\mathbf{a}$ (which takes binary value $a_{n}\in\{0,1\}$, where 0 reflects not allocated, or 1 reflects allocated) and the control impact vector $\mathbf{c_{i}}$ towards objective $i$, containing the impact of control variables $c_{1},c_{2},...,c_{N}$ directly related towards control variables $x_{1},x_{2},...,x_{N}$.

### v2

@@ Abstract
Strategic project planning and dynamic control are essential to ensure that complex projects are executed in a best-fit-for-common-purpose manner. When such strategies become inconceivable for humans, effective and efficient computer-aided decision support becomes crucial. Standard simulation-driven evaluation and a-posteriori decision-making are often used, rather than a combined simulation and optimisation approach that a-priori integrates the stakeholder interests and risk management. However, current planning and control methodologies often lack stochastic representation and/or associative multi-objective optimisation approaches within socio-technical reach. This paper introduces Odycon (Open design and dynamic control), a new purpose-driven project management by design methodology that resolves these shortcomings. A generic mathematical project planning and control statement is developed to integrate and represent both stakeholders preferences (human domain) and project system performances (technical domain) for effective project planning and dynamic control. Odycon integrates Monte-Carlo simulation (MCS) with the Integrative Maximisation of Aggregated Preferences (IMAP) optimisation method to develop a best-fit strategic plan and dynamic control measures. Its value is demonstrated through applications in offshore wind installation and a highway infrastructure projects, showcasing advances in associative design and decision-making, and aiming for a best-fit-for-common purpose synthesis across different complex project phases.
@@
Project management , Project planning & scheduling , Stochastic Project control , Risk management , Multi-objective optimisation , Preference & Performance-based optimization , Monte Carlo simulation , Preference function modelling
@@
[…]
[c]organization=Department Algorithmics, Faculty of Mathematics and Computer Science, Delft University of Technology,addressline=Van Mourik Broekmanweg 6, city=Delft, postcode=2628 XE, state=Zuid-Holland, country=The Netherlands
@@ 1 Introduction
In today’s complex engineering project management landscape, wicked problems often cause projects to derail. In fact, projects don’t go wrong — they start wrong (Flyvbjerg, 2024). And even when they start right, they can still lose direction. Advanced design and project management practices may still hit impasses because project managers often rely on retrospective analyses (such as those by Flyvbjerg (2024) or others like Yang et al. (2022)) and traditional decision-support methods, which offer little guidance for a constructive way forward. These challenges become particularly acute when decision-makers fail to recognize that their problems are part of a larger, interconnected whole — and that mutual interconnection and collaboration are key to unlocking complex projects that achieve a best-fit for a common purpose.
This calls for moving beyond the traditional notion of fitness for purpose (total quality), which emphasizes meeting predefined requirements and delivering technically effective results. Instead, we must adopt a broader, value-driven perspective — one that focuses on what is socially desirable and physically feasible from a shared purpose. Best-fit for common purpose integrates stakeholder interests, dynamic context, and collective value creation within socio-physical reach.
Moreover, what is needed is a proactive threefold approach to confront project complexity — ensuring that projects start, proceed, and finish successfully: (1) Agreeing First on a best-fit scope, (2) Acting Feasibly with an optimal project plan, and (3) Adapting Flexibly through the best set of mitigation measures. Establishing a robust and socio-technically feasible project plan is fundamental to effective and efficient project management (Smith et al., 2014). However, as project size and uncertainty increase, initial estimates regarding scope, cost, and timelines are often exposed to inherent risks and uncertainties that cannot be entirely eliminated.
Project success is therefore achieved through dynamic adaptation to emerging challenges—often not specified in a static contract—and through collaborative engagement among stakeholders to develop associative solutions. Focusing solely on the final scope as defined in the contract, insisting on creating a fixed plan and adhering rigidly to it, is ultimately naïve and pseudo-realistic in the face of real-world project dynamics.
This ultimately calls for an actionable, open design and systems-oriented approach, supported by transparent decision-support tools that foster both deliberate (slow) and socially informed thinking. As such, successful project delivery hinges on purpose-driven and optimal strategic planning and dynamic control, transforming complexity into outcomes that best fit within socio-physical reach. Planning refers to the process of reaching agreement first in order to act feasibly, while project control ensures execution according to plan and the ability to adapt flexibly when needed on the run.
Despite this reality, traditional project management in construction often remains heavily focused on cost and schedule, with stakeholders primarily considering these from individual perspectives. This narrow focus neglects the interplay of stakeholder-oriented, concurrent objectives. Joint project success can only be enhanced when stakeholders are willing to look beyond self-interest and engage in a process of mutual alignment. This requires stakeholders to let go of their individually defined objectives in favour of a synthesis solution that best fits the common purpose — one that ultimately outperforms single-sided or even compromise-based project outcomes (Wolfert, 2023; Scharmer, 2016; Glasl, 1998; Van den Doel and Van Velthoven, 1993).
By integrating diverse stakeholder interests and accounting for project capabilities in the decision-making process, a more holistic outcome—referred to as the best-fit for common purpose—can be achieved, rather than settling for suboptimal compromises (Zhilyaev et al., 2022). This best-fit design emerges from the socio-technical interplay between stakeholder preferences (the human domain—what they want) and system performance capabilities (the technical domain—what it can deliver), as developed and demonstrated by Van Heukelum et al. (2024).
However, achieving such a best-fit requires more than technical alignment; it also demands a fundamental shift in how planning and execution control are approached. Both academic research and industry practice tend to rely on retrospective, analysis-oriented methods (a-posteriori), which offer limited support for forward-looking strategic development. Moreover, these methods are often driven by individual optimization or simulation-based approaches, producing one-sided solutions that overlook the multifaceted nature of real-world projects. In practice, technical capabilities, human goal-oriented behavior, and stakeholder desirability are deeply interconnected. Yet traditional planning and control approaches fail to explicitly capture the dynamic interplay of all stakeholder interests within this complex decision-making process (Blanchard and Fabrycky, 2011; Del Pico, 2023; Slack et al., 2010).
Most existing design optimisation methods rely on single-objective optimisation (SOO), often reducing multiple project objectives to a single cost metric. This approach overlooks the qualitative nature of stakeholder preferences, which are essential for goal-oriented decision-making. While the benefits of multi-objective optimisation (MOO) in project management are increasingly recognized (Guo and Zhang, 2022), current methodologies rarely integrate MOO within strategic project planning or dynamic project control frameworks that also incorporate probabilistic risk management (Kammouh et al., 2022). To address the rising complexity and uncertainty in construction projects, decision support must evolve to enable computer-aided, stakeholder-oriented planning and control based on MOO. As Kahneman (2011) emphasizes, deliberate and data-informed processes are essential for mitigating cognitive biases such as overconfidence. Consequently, well-informed project decision-making requires transparent, evidence-based tools that support “slow and social thinking”—tools that not only process data but also help align diverse stakeholder objectives within feasible and resilient project solutions.
In this paper, all the aforementioned considerations are taken into account in the development of a new open-design-driven dynamic project planning and control methodology, named Odycon (acronym: Open Design and Dynamic Control). Odycon builds on the core principles and optimization method of the Open Design Systems (Odesys) methodology (Wolfert, 2023). The Odesys integrative maximisation of aggregated preferences (IMAP) optimization method has proven effective in achieving best-fit-for-common-purpose design solutions, where all stakeholder interests are accounted for and translated into a common preference domain to identify an overall group optimum. Notably, Odesys/IMAP is a novel preference-based optimization methodology where, instead of directly minimizing or maximizing individual objectives, the rigorous mathematical optimization focuses solely on maximizing the aggregated preference of stakeholder objectives within physical reach.
So far, the Odesys/IMAP method enables iterative group design and decision-making in a deterministic manner. However, for a project planning, it must be extended to incorporate traditional probabilistic project planning techniques that capture schedule performance, such as PERT or discrete-event simulation (DES). This constitutes a truly novel project management-by-design methodology. Additionally, these tools should employ stochastic simulation methods to reflect the inherent uncertainties and unpredictability encountered both in the planning and execution phases. For project control applications he method must also be expanded to shift the focus from optimizing design variables to optimizing the allocation of control measures, using multiple project objectives to evaluate their effectiveness and efficiency. To achieve this, the state-of-the-art project manager-oriented mitigation controller (MitC), as developed and demonstrated by Kammouh et al. (2022, 2021), will be enhanced from a single-cost objective orientation to optimally facilitate stochastic, multi-objective project control. In doing so, Odycon constitutes a truly novel project management-by-design methodology.
To summarize, the following Odycon advances are presented in this paper: 1. A general framework for integrated strategic planning and dynamic control, combining probabilistic project schedule simulations with preference-based multi-objective design optimisation. This enables dynamic adaptation to project-specific conditions and stakeholder objectives, thereby enhancing associative decision-making in complex construction environments. 2. A direct linkage between project feasibility, project management objectives, and stakeholder decisions, resulting in a unified integration of network performance, stakeholder values, and associative planning and control processes. 3. A pure preference-based aggregation mechanism for multi-objective decision-making in project management. This translates diverse objectives and constraints into a shared preference domain, enabling an optimised group synthesis without relying on monetisation. Rather than selecting the lowest cost solution, the approach supports identifying the solution with the highest combined value for both project outcomes and people—that is, a best-fit for common purpose plan and set of control measures. 4. A methodological advancement from PII to PIII in project management practice. Building on the PII model for complex construction projects introduced by Van Gunsteren (2011), itself rooted in double-loop learning as described by Schön and Argyris (1996), this paper proposes an evolution toward PIII—Best Practice for Collaborative Projects. Unlike PII, which focuses on making improved decisions for internal stakeholders via integrative reflection and learning, PIII promotes associative open-loop learning and co-creation with all relevant stakeholders. This fosters dynamic group-based planning and control, where integrative maximisation of associated preferences replaces static, self-oriented decision-making. In doing so, it enables a genuinely collaborative approach to achieving the best possible overall outcome for both project and people (Wolfert, 2023).
Combining the advances of MitC and IMAP enables a comprehensive representation of all stakeholder preferences and objectives in project planning and control, while simultaneously accounting for socio-physical constraints. This significantly reduces the risk of inefficient or misinformed decision-making—especially as project size and complexity increase, making it virtually impossible to manually consider all potential scenarios. Odycon supports the generation of best-fit-for-common-purpose solutions for both strategic planning and dynamic control applications. As a decision-support methodology, Odycon automates the selection of planning and control variables by integrating multiple stakeholder objectives and constraints into a unified stochastic simulation and optimisation process.
@@
This paper is structured as follows. It begins by introducing the general mathematical formulation of Odycon’s multi-objective project planning and control optimisation approach. Next, the underlying framework and core concepts of the stochastic simulation and optimisation methodology are explained, including a detailed description of the method’s design, within which the IMAP optimisation is embedded. The methodology is first demonstrated using a strategic planning case (with no control variables) involving an offshore wind installation project. In this scenario, a service provider and a marine contractor act as the primary decision-makers and are able to agree on a shared scope, resulting in an actionable and feasible optimal plan. Following this, a dynamic control case (with no planning variables) is presented, applied to a highway infrastructure project. This demonstrator features associative mitigation control by both a contractor and a highway agency, illustrating how to arrive at a best set of control measures to enable flexible adaptation across two distinct project management phases. The paper concludes with recommendations for further development and reflects on the added value of the newly developed stochastic simulation and optimisation methodology.
@@ 2 Mathematical formulation of the multi-objective optimisation
    $P_{k,i}(O_{i}(\mathcal{N}(\mathbf{x},\mathbf{y}),F_{1,2,...,J}(\mathbf{x},\mathbf{y}))$: Preference functions that describe the preference stakeholder $k$ has towards objective function $i$, which are functions of different performance functions and dependent on planning and/or control variables and (physical) variables ( $i\leq k$ and $K$ is the maximum number of stakeholders).
@@
    $\mathcal{N}(\mathbf{x},\mathbf{y})$: Performance function that describes the project network, depending on planning and/or control variables $\mathbf{x}$ and (physical) variables $\mathbf{y}$. Here $\mathcal{N}$ is a logical representation of a project planning network, composed of nodes (activities) and edges (logical links) with their respective properties.
@@
    $\mathcal{N}(\mathbf{x},\mathbf{y})$: Performance function that describes the project network, depending on planning and/or control variables $\mathbf{x}$ and physical variables $\mathbf{y}$. Here, $\mathcal{N}$ is a parametric graph representation of a project planning network, composed of nodes (activities) $V$ and edges (logical links) $E(\mathbf{x},\mathbf{y})$. Each node $v\in V$ has attributes such as duration $D_{v}(\mathbf{x},\mathbf{y})$, risk $R_{v}(\mathbf{x},\mathbf{y})$, and cost $C_{v}(\mathbf{x},\mathbf{y})$, while each edge $e\in E(\mathbf{x},\mathbf{y})$ has attributes such as weight $W_{e}(\mathbf{x},\mathbf{y})$ and capacity $C_{e}(\mathbf{x},\mathbf{y})$.
  • 7.
@@
The preference aggregation is performed using the principles of the PFM theory (Barzilai, 2022). It is an integral part of the Integrative Preference Aggregation Method (IMAP), see Wolfert (2023). A best-fit solution is identified as a feasible plan that yields the highest aggregated preference. Note that individual preferences reflect the level of interest in specific objectives. To arrive at an aggregated preference for individual objectives, the IMAP method employs the mathematical operator, A, which determines the aggregated preference score using an algorithm called the A-fine Aggregator (see A).
@@
The preference functions describe the relation of a stakeholders satisfaction towards a certain objective on a defined scale. In other words, preference — the only property of relevance in decision theory and equivalent to value — provides goal orientation to the objective function. Note: objective functions, in themselves, merely describe an objective and do not inherently imply a minimisation or maximisation direction — a distinction that sets IMAP apart from all classical multi-objective optimisation methods. The elicitation of the preference functions and associated individual weights is essential to reflect the relationship of the stakeholders interest towards a certain objective. As an initial estimation for both the preference functions input and local and global the weights the Conjoint Analysis (CA) method can be used. For more information regarding the process of preference elicitation, the reader is referred to Arkesteijn et al. (2017). As mentioned in Van Heukelum et al. (2024) objective $O_{i}$ can be associated with multiple stakeholder preference functions $P_{k,i}$ (as $k\geq i$). Therefore, it is not required that a stakeholder expresses a preference for all objectives.
@@
In the context of a pure project planning optimisation (no control measures), the variables $x_{n}$ are defined as stated above, allowing to take any form (e.g. integer, float, binary, etc.). However, for a pure project control optimisation (no planning variables), the variables $\mathbf{x}$ are redefined as $\mathbf{x}=\mathbf{a}\mathbf{c_{i}}$, that is the product of the allocation of a control variables vector $\mathbf{a}$ (which takes binary value $a_{n}\in\{0,1\}$, where 0 reflects not allocated, or 1 reflects allocated) and the control impact vector $\mathbf{c_{i}}$ towards objective $i$, containing the impact of control variables $c_{1},c_{2},...,c_{N}$ directly related towards control variables $x_{1},x_{2},...,x_{N}$. Note that $x_{1..N}\in\{0,c_{1..N}\}$, which implies that the solution space is bounded.